OpenStudy (anonymous):

Subtract six and three-eighths minus negative two and seven-eighths. nine and one-fourth three and one-half seven and one-fourth two and five twenty-thirds

OpenStudy (texaschic101):

6 3/8 - (-2 7/8) = (change to improper fractions) 51/8 - (-23/8) = 51/8 + 23/8 = 74/8 which equals 9 1/4